# Graph Residual based Method for Molecular Property Prediction

### v1

@@ Abstract
Property prediction of materials has recently been of high interest in the recent years in the field of material science. Various Physics-based and Machine Learning models have already been developed, that can give good results. However, they are not accurate enough and are inadequate for critical applications. The traditional machine learning models try to predict properties based on the features extracted from the molecules, which are not easily available most of the time. In this paper, a recently developed novel Deep Learning method, the Graph Neural Network (GNN), has been applied, allowing us to predict properties directly only the Graph-based structures of the molecules. SMILES (Simplified Molecular Input Line Entry System) representation of the molecules has been used in the present study as input data format, which has been further converted into a graph database, which constitutes the training data. This article highlights the detailed description of the novel GRU-based methodology to map the inputs that have been used. Emphasis on highlighting both the regressive property as well as the classification-based property of the GNN backbone. A detailed description of the Variational Autoencoder (VAE) and the end-to-end learning method has been given to highlight the multi-class multi-label property prediction of the backbone. The results have been compared with standard benchmark datasets as well as some newly developed datasets. All performance metrics which have been used have been clearly defined as well as their reason for choice.
@@
GNN , VAE, SMILES, multi-label multi-class classification , GRU
@@ 1 Introduction
Quantitative Structure Property Relationship (QSPR) a traditional method developed as early as 1935 [1], has provided an efficient framework to facilitate computation on chemical properties by breaking them down into their smallest blocks. The mathematical model allows the molecules to be broken down according to their chemical, structural, physical, and biological properties. The method of property prediction consists of two steps which involve breaking down the molecule into property descriptors followed by regression on these molecules. The molecular descriptors generate a n-dimensional data with $d=[d_{1},d_{2},\ldots,d_{n}]^{T}$ which indicates the presence of a particular feature as generated by the molecular descriptor.
@@
The molecular descriptors can be 0D,1D,2D as well as 3D as described in the [2, 3]. The most common type of descriptors is 0D descriptors which is typically used in ML models as used in Asheri et al [4] work. The 0D molecular descriptors are simply atoms and molecular properties that are used to describe the molecule as defined by [5]. However, these descriptors don’t contain any information of the structural or connectivity between atoms. Several works can be found using these 0D methods [6]. 1D descriptors have been mostly used till now, they can be thought of as functional sub-groups or the building blocks in a molecule. Most of the works that have been done till now are using this method, however, since these don’t contain any structural information, their accuracy is not up to par. In more recent times, the use of 2D descriptors has been used, which allows the molecular topology to be represented in its true form and introduces several geometry-based features that can be represented in graph format like cyclicity, symmetry, and connectivity between atoms. Lastly, a more modern higher dimensional descriptors [7] which can represent complex molecular geometries and dynamics has been used in cases of [8, 9, 10] however, exploiting such features is computationally expensive.
@@
QSPR modeling is a challenging process as it requires subject expertise to identify the functional subgroups. Automating the process, it requires a careful selection of algorithms and data to identify the information descriptor that is to be used in the ML model. This inherently poses a significant challenge since such descriptors cannot be generalized for all tasks and can be found as trial and error. Due to all the above reasons, alternate ways have come up to extract molecular embeddings or fingerprints directly from the molecular structure rather than from its molecular descriptors. This is where Graph Neural Networks (GNNs) offer efficient methodology. A molecule can be arranged in a graph-like structure from which the deep learning module will directly extract the features. This kind of learning on non-Euclidean datasets [50] is also known as geometric-learning and has been made more easily possible due to the recent development of Pytorch-geometric [11].
GNNs are recognized as a potent tool for extracting knowledge from graph-structured data, showcasing impressive performance across a variety of domains. These domains include social networks, molecular structures, drug screening, chip design, traffic flows, recommendation systems, and knowledge graphs. In various tasks such as node classification, graph classification, graph regression, and link prediction, GNNs consistently achieve state-of-the-art performance. The concept of GNN was first introduced by the paper [12] where they extended the recursive neural networks to give rise to a new set of networks that is more node-focused. The work was further extended by Scarselli et al [13] which gave rise to an end-to-end network of GNN along with modern backpropagation techniques. In this work, the authors provided a detailed mathematical description of the GNN method. One of the major important properties of Graphs and GNNs are that they are size-agnostic i.e., the sample size can differ from one input to another as well as permutation invariant. This property provides a major advantage to GNNs over CNN-based architectures since the size of the input graph i.e. the number of nodes can easily vary whereas in CNNs we must provide a fixed set of input size.
There are various tasks that can be performed by GNNs such as node level tasks as mentioned in [14, 15], edge level tasks such as [16] and Graph level tasks as described in [17, 18]. The problem of Molecule property prediction which we are trying to solve is a Graph level task where each molecular graph generates an embedding which is to be further used. The most primitive GNN used for solving such type of problem has been in the form of Message Passing Neural Network (MPNN) [19]. The MPNN network is based on the simple Graph Convolution format which allows each node to learn from its subsequent neighbouring nodes and that information is then aggregated. Similar, to the Graph Convolutions that are used in MPNN, several other aggregators have been proposed, most notable being Graph Attention Network (GAT) based on the multi-head attention mechanism [20], the recently developed GraphSAGE (Graph Sample and Aggregated) [21], a framework for inductive representation learning on large graphs which samples and aggregates feature from a node’s neighbourhood to learn its embedding, GraphConv mechanism [22] etc. Graph Isomorphism Network (GIN) by Xu eta al [23] provides flexibility in capturing the local and global graph structure, making it suitable for regression tasks.
This paper emphasizes predicting both the regression and classification properties of molecules. It presents results for new datasets and compares them with well-established benchmarks. The classification section specifically examines organic and inorganic compounds, highlighting our research’s effectiveness in drug discovery and predicting molecular properties. Graph classification holds particular significance in the realm of molecular property prediction, where molecular structures are represented as graphs, with atoms as nodes and chemical bonds as edges. In this context, graph classification tasks play a crucial role in predicting various molecular properties, such as toxicity, solubility, bioactivity, and drug efficacy, as mentioned in [24, 25] among others. By leveraging graph-based representations, researchers can capture intricate structural features and relationships inherent in molecular compounds, enabling more accurate and comprehensive predictions. Tasks within molecular property prediction often include binary classification to distinguish between toxic and non-toxic compounds, multiclass classification to categorize compounds into different therapeutic classes or activity levels, and regression to predict quantitative properties like molecular weight or solubility. These tasks are pivotal in drug discovery, materials science, and environmental health.
Jiying et al [26] presented an approach to fine-tuning Graph Neural Networks by framing the problem within an Optimal Transport framework, capitalizing on graph topology. Through the introduction of a masked Optimal Transport problem and the definition of the Masked Wasserstein distance, their GTOT-Tuning (Graph Induced Optimal Transport) framework facilitates efficient knowledge transfer from pre-trained models while preserving local feature invariances. They leverage graph structure to enhance knowledge transfer and offer flexibility for integration with complementary techniques like contrastive learning [27] . However, this method poses a high computational demand and has limited efficacy on small datasets or those with sparse graph structures, as is usually the case with medicinal and molecular data.
Another enhancement on GNNs is the Graph Multiset Transformer (GMT), introduced by Jinheon et al [28] as a graph multi-head attention-based global pooling layer. It captures interactions and relations between nodes of a graph based on its structural dependencies. This ensures that GMT satisfies both injectiveness and permutation invariance. While leading the model to learn more accurate representations of entire graphs by capturing structural dependencies between nodes, training a transformer model such as GMT requires significant computational resources. The inner workings of GMT also capture complex structural dependencies, which are hard to interpret in a physical manner.
Another important context that has been explored is the context of multi-class multi-label classification. Multilabel classification is widely explored in the Image modality, with extensive use of transformer-like encoder-decoder models, as shown by Shichao et al [29] and Shilong et al [30]. Graphical Classification relies heavily on Transformer based approaches as well, with GNN based models lacking severely in performance.
Various methods exist for multi-class multi-label methods on different modalities of input data such as [31, 32]. In this context, we are trying to explore the classification capabilities of the GNNs. The first paper proposed this was a simple GCN structure [33] which was used for multi-label classification.
With the introduction of “Attention is all you need” paper [34] and recent advances in Transformer architecture, Shermukhamedov et al [35] introduced ElemBert – v1 and ElemBert – v2, which employs a transformer mechanism for binary classification utilizing structural information. Their models, known for their flexibility, demonstrate adaptability across various datasets. Through benchmarking, they have showcased the state-of-the-art performance of their models in diverse material property prediction tasks, encompassing both organic and inorganic compounds. Additionally, Heyrati et al [24] proposed BioAct-Het, leveraging a heterogeneous Siamese neural network to capture the intricate relationship between drugs and bioactivity classes, thereby consolidating them within a unified latent space. While this approach enhances the identification of drug discovery markers, it faces limitations in effectively handling the simultaneous classification of multiple labels. Notably, both approaches require separate models for training each label, resulting in heightened training and inference times, as well as increased resource requirements per data point. Alpersein et al [36] introduce a new method called the All SMILES Variational Autoencoder (VAE), focusing on enhancing the optimization of molecular properties through VAEs that work with both SMILES string and graph-based molecular representations. This approach [36] handles the encoding of a molecule’s multiple SMILES strings through a layered arrangement of recurrent neural networks. It aggregates the hidden states corresponding to each atom across these representations, employing an attention mechanism to consolidate them into a comprehensive, fixed-size latent vector. The model’s ability to decode this representation back into a distinct set of SMILES strings for the original molecule allows it to establish a nearly one-to-one relationship between the molecule and its latent representation. This technique effectively addresses the issue of SMILES strings’ non-uniqueness and sidesteps the computational demands associated with graph convolution operations. The utilization of VAEs is particularly advantageous in this context because they are adept at learning intertwined representations of complex data distributions. Additionally, VAEs incorporate a regularization mechanism by enforcing a prior distribution on the latent space, which aids in mitigating overfitting and enhances the model’s ability to generalize.
In this paper, an attempt has been made to introduce a novel architecture GRU-based Residual ECC network to capture the inherent pattern in the molecular graphs and introduce long-range dependencies. A complete model overview is presented along with its mathematical formulation to explain the underlying architecture. Apart from that, a diagrammatic representation of the model in an end-to-end fashion has been presented for a better understanding of the workflow for both regression and classification tasks. The highlight of the work is the novel GRU-based Residual GNN, along with its explainability and how it relates to the underlying chemistry has been presented. In this work, a comparison has been made against the SOTA model, which is present, and the performance on benchmark datasets. A detailed statistical analysis of the classification datasets as well as the data processing, has also been described in this study.
@@ 2 Graphical Representation of Molecule
Morgan’s technique [37] for generating unique machine descriptions and the establishment of the CAS ONLINE search system were significant developments for the description of the chemical structure database. The widespread adoption of computers in the field of chemical nomenclature led to the prevalent use of line notation. Instead of drawing detailed molecular structures, line notation uses lines, symbols, and characters to convey essential structural information. This is attributed to the computational efficiency with which computers can handle linear data strings [38]. Rapid advancements in computer technology have outpaced the incremental growth of chemical information. The synergy between computer capabilities and chemical knowledge allows for the storage of all existing chemical information on current hardware. The historical challenge of merely storing information in machines has been largely overcome. Current and future efforts are now focused on creating highly efficient systems that deliver chemically relevant information when required. To achieve this, a new chemical language and associated computer programs are being developed, grounded in the innovative information system known as SMILES [39].
@@
SMILES is a line notation used to encode molecular structures, allowing for a linear representation of the molecule so that a computer program can understand the structure. It specifically represents a valence model of a molecule and is widely used in chemistry for representing molecular structures in a compact and human-readable form. The graph of a chemical structure was to be uniquely described, including but not limited to the molecular graph comprising nodes (atoms) and edges (bond) [40]. Since most of the datasets, comes in SMILES [39] encoded format, it needs to be converted into graph database for use in the GNN. A molecular graph can be represented in the form of $G(n)=(V,E)$ where, $n$ is the number of nodes or vertices i.e the atoms in the molecule and $e_{v,w}$ represents the edges i.e the atomic bonds in the molecule where $e_{v,w}\epsilon E$. A typical example of the molecule in this format has been shown in the Figure 1.
@@
Atom Type Atomic Number Neighbour Hydrogen Formal Charge
@@
Table 2: Representation of encoding of edge features
Bond Conjugated Aromatic Is In Ring Connects
Edge 1 Single No No No (1,2)
Edge 2 Double No No No (2,3)
Edge 3 Single No No No (2,4)
Table 3: Node features
Atom Type All possible atom list in the periodic table 117 [0-116]
Bonds Number of bonds the atom is involved in 9 [117-125]
Charge Formal Charge of the atom 1 [126]
Hybridization sp, sp2, sp3, sp3d, sp3d2, sp2d, s, other, unspecified 9 [127-135]
Is Aromatic Indicate if the atom forms any aromatic system 1 [136]
Is in Ring Indicate if the atom forms any ring 1 [137]
Atomic Number Atomic Number 1 [138]
Hs Number of bonded hydrogen atoms 9 [139-147]
Table 4: Edge features
@@
Bond Type Single, Double, Triple and Aromatic 4 [0-3]
@@
There are several types of atomic as well as bond features present, however, for our study, we use only the basic features that will be available without experimentation to allow for the generalizability of the model. Based on Figure 1 a sample encoding structure of the features are presented in the Tables 1 and 2 along with their encodings. These numerical encodings are then converted to one-hot encoded vectors for use in the geometric dataset. The tables show a small sample of the node and edge feature set.
A complete representation of the number of number of features used for both the node and edge cases respectively, has been shown in Tables 3 and 4. They show the size of the feature space of the one-hot encoded vectors. Proper dimension size along with the type of features with detailed descriptions are presented here.

### v2

@@ Abstract
Machine learning-driven methods for property prediction have been of deep interest. However, much work remains to be done to improve the generalization ability, accuracy, and inference time for critical applications. The traditional machine learning models predict properties based on the features extracted from the molecules, which are often not easily available. In this work, a novel Deep Learning method, the Edge Conditioned Residual Graph Neural Network (ECRGNN), has been applied, allowing us to predict properties directly only the Graph-based structures of the molecules. SMILES (Simplified Molecular Input Line Entry System) representation of the molecules has been used in the present study as input data format, which has been further converted into a graph database, which constitutes the training data. This manuscript highlights a detailed description of the novel GRU-based methodology, ECRGNN, to map the inputs that have been used. Emphasis is placed on highlighting both the regressive property and the classification efficacy of the same. A detailed description of the Variational Autoencoder (VAE) and the end-to-end learning method used for multi-class multi-label property prediction has been provided as well. The results have been compared with standard benchmark datasets as well as some newly developed datasets. All performance metrics that have been used have been clearly defined, and their reason for choice.
@@
ECRGNN,VAE,SMILES,multi-labelmulti-class classification,GRU
IIT Bombay] Department of Mechanical Engineering, IIT Bombay, Mumbai, India CMINDS, IIT Bombay] Centre for Machine Intelligence and Data Science, IIT Bombay, Mumbai, India IIT Bombay] Department of Mechanical Engineering, IIT Bombay, Mumbai, India IIT Bombay] Department of Mechanical Engineering, IIT Bombay, Mumbai, India
{tocentry}[Uncaptioned image]
Overall Pipeline for Molecular Property Prediction
@@ 1 Introduction
Quantitative structure-property relationship (QSPR) a traditional method developed as early as 1935 1, has provided an efficient framework to facilitate computation on chemical properties by breaking them down into their smallest blocks. The mathematical model allows the molecules to be broken down according to their chemical, structural, physical, and biological properties. The method of property prediction consists of two steps, which involve breaking down the molecule into property descriptors and then regressing on these descriptors. The molecular descriptors generate a n-dimensional data with $d=[d_{1},d_{2},\ldots,d_{n}]^{T}$ which indicates the presence of a particular feature as generated by the molecular descriptor.
@@
Molecular descriptors can be 0D, 1D, 2D, as well as 3D. 2, 3 The most common type of descriptors is 0D descriptors, typically used in conjunction with Machine Learning models. 4 The 0D molecular descriptors are simply atoms and molecular properties that are used to describe the molecule. 5 However, these descriptors do not contain any information about the relationship between the atoms. 1D descriptors can be considered functional sub-groups or a molecule’s building blocks. It is the most commonly used descriptor, yet these don’t contain any structural information. Thus, most of the work that has been done using this method to this day lacks accuracy. In more recent times, the use of 2D descriptors has been used, which allows the molecular topology to be represented in its true form and introduces several geometry-based features that can be represented in graph format, like cyclicity, symmetry, and connectivity between atoms. Lastly, a more modern higher dimensional descriptors 6, which can represent complex molecular geometries and dynamics, has been used 7, 8, 9; however, exploiting such features is computationally expensive.
@@
QSPR modelling is challenging as it requires subject expertise to identify the functional subgroups. Automating the process requires a careful selection of algorithms and data to identify the information descriptor that is to be used in the Machine Learning model. This inherently poses a significant challenge since such descriptors cannot be generalized for all tasks. Thus, alternate ways have come up to extract molecular embeddings or fingerprints directly from the molecular structure rather than from its molecular descriptors. This is where Graph Neural Networks (GNNs) offer efficient methodology. A molecule can be arranged in a graph-like structure from which deep learning modules directly extract relevant features. This kind of learning on non-Euclidean datasets is also known as geometric learning and has been made more easily possible due to the recent development of the Pytorch-geometric module. 10
@@ 2 Graphical Representation of Molecule
GNNs are recognized as a potent tool for extracting knowledge from graph-structured data, showcasing impressive performance across a variety of domains, like social networks, molecular structures, drug screening, chip design, traffic flows, recommendation systems, and knowledge graphs. In various tasks such as node classification, graph classification, graph regression, and link prediction, GNNs consistently achieve state-of-the-art performance. The concept of GNN was first introduced by Gori et al. 11, where they extended the recursive neural networks to give rise to a new set of networks that is more node-focused. This work was further extended by Scarselli et al. 12, which created an end-to-end GNN network along with modern backpropagation techniques. The authors provided a detailed mathematical description of the GNN method in this work. One of the major important properties of graphs and GNNs is that they are size-agnostic, i.e., the sample size can differ from one input to another, and they are permutation invariant. This property provides a major advantage to GNNs over CNN-based architectures since the input graph size and the number of nodes and edges can easily vary, whereas, in CNNs, we must provide a fixed set of input sizes.
There are various variety of tasks that can be performed by GNNs, such as node level tasks 13, 14, edge level tasks 15, and Graph level tasks 16, 17. The problem of Molecule property prediction, which we are trying to solve, is a Graph-level task where each molecular graph generates an embedding, which is to be used further for prediction. The most primitive GNN used for solving such types of problems is the Message Passing Neural Network (MPNN). 18 MPNN is based on the simple Graph Convolution format, which allows each node to learn from its subsequent neighbouring nodes, and that information is then aggregated into the original node. Similar to the Graph Convolutions that are used in MPNN, several other aggregators have been proposed, most notable being Graph Attention Network (GAT) based on the multi-head attention mechanism 19, the recently developed GraphSAGE (Graph Sample and Aggregated) 20, a framework for inductive representation learning on large graphs which samples and aggregates feature from a node’s neighbourhood to learn its embedding, the GraphConv mechanism 21 etc. Graph Isomorphism Network (GIN) by Xu et al. 22 provides flexibility in capturing the local and global graph structure, making it suitable for regression tasks.
Jiying et al. 23 presented an approach to fine-tune Graph Neural Networks by framing the problem within an Optimal Transport framework, capitalizing on graph topology. By introducing a masked Optimal Transport problem and defining the Masked Wasserstein distance, the Graph Induced Optimal Transport (GTOT) -Tuning framework facilitates efficient knowledge transfer from pre-trained models while preserving local feature invariances. GTOT leverages graph structure to enhance knowledge transfer and offer flexibility for integration with complementary techniques like contrastive learning. 24 However, this method poses a high computational demand and has limited efficacy on small datasets or those with sparse graph structures, as is usually the case with medicinal and molecular data.
Another enhancement on GNNs is the Graph Multiset Transformer (GMT), introduced by Jinheon et al. 25 as a graph multi-head attention-based global pooling layer. It captures interactions and relations between nodes of a graph based on its structural dependencies. This ensures that GMT satisfies both injectiveness and permutation invariance. While this leads the model to learn more accurate representations of entire graphs by capturing structural dependencies between nodes, training a transformer model such as GMT requires significant computational resources. The inner workings of GMT also capture complex structural dependencies, which are hard to interpret in a physical manner.
Multilabel classification is widely explored in the image modality, with extensive use of transformer-like encoder-decoder models, as shown by Shichao et al. 26 and Shilong et al. 27. Graphical classification also relies heavily on transformer-based approaches, with GNN-based models severely lacking performance. Various methods exist for multi-class multi-label methods on different modalities of input data. 28, 29 In this context, we are trying to explore the classification capabilities of GNNs. The first paper that proposed this was a simple Graph Convolutional Network (GCN) structure 30, which was used for multi-label classification.
With the introduction of the "Attention is all you need" paper 31 and recent advances in Transformer architecture, Shermukhamedov et al. 32 introduced ElemBert – v1 and ElemBert – v2, which employ transformer mechanisms for binary classification utilizing structural information. These models, known for their flexibility, demonstrate adaptability across various datasets. Through benchmarking, they have showcased the state-of-the-art performance of Elembert in diverse material property prediction tasks, encompassing both organic and inorganic compounds. Additionally, Heyrati et al. 33 proposed BioAct-Het, leveraging a heterogeneous Siamese neural network to capture the intricate relationship between drugs and bioactivity classes, thereby consolidating them within a unified latent space. While this approach enhances the identification of drug discovery markers, it faces limitations in effectively handling the simultaneous classification of multiple labels. Notably, both approaches require separate models for training each label, resulting in heightened training and inference times and increased resource requirements per data point.
Alpersein et al. 34 introduce a new method called the All SMILES Variational Autoencoder (VAE), focusing on enhancing the optimization of molecular properties through VAEs that work with both SMILES string and graph-based molecular representations. This approach handles encoding a molecule’s multiple SMILES strings through a layered arrangement of recurrent neural networks. 34 It aggregates the hidden states corresponding to each atom across these representations, employing an attention mechanism to consolidate them into a comprehensive, fixed-size latent vector. The model’s ability to decode this representation back into a distinct set of SMILES strings for the original molecule allows it to establish a nearly one-to-one relationship between the molecule and its latent representation. This effectively addresses the issue of SMILES strings’ non-uniqueness and sidesteps the computational demands associated with graph convolution operations. VAEs are particularly advantageous in this context because they are adept at learning intertwined representations of complex data distributions. Additionally, VAEs incorporate a regularization mechanism by enforcing a prior distribution on the latent space, which aids in mitigating overfitting and enhances the model’s ability to generalize.
This work emphasizes the prediction of both the regression and classification properties of molecules. It presents results for new datasets and compares them with well-established benchmarks. The implemented Graph-based classification shows significance in the realm of molecular property prediction, where molecular structures are represented as graphs, with atoms as nodes and chemical bonds as edges, playing a crucial role in predicting various molecular properties such as toxicity, solubility, bioactivity, and drug efficacy, among others. 33, 35 Intricate structural features and relationships inherent in molecular compounds can be captured by leveraging graph-based representations, enabling more accurate and comprehensive predictions. Tasks within molecular property prediction often include binary classification to distinguish between toxic and non-toxic compounds, multiclass classification to categorize compounds into different therapeutic classes or activity levels, and regression to predict quantitative properties like molecular weight or solubility. These tasks are pivotal in drug discovery, materials science, and environmental health.
This paper attempts to introduce a novel architecture, a Gated Recurrent Unit (GRU) - based residual edge convolutional network to capture the inherent pattern in the molecular graphs and introduce long-range dependencies. A complete model overview is presented along with its mathematical formulation to explain the underlying architecture. Apart from that, a diagrammatic representation of the model in an end-to-end fashion has been presented to better understand the workflow for both regression and classification tasks. This study’s highlight, along with its explainability in relation to the underlying chemistry. Additionally, a comparison has been made between the present SOTA model and the performance of benchmark datasets. This study has also described a detailed statistical analysis of the classification datasets and data processing.
SMILES Encoding
@@
Morgan’s technique 36 for generating unique machine descriptions and the establishment of the CAS ONLINE search system were significant developments for the description of the chemical structure database. The widespread adoption of computers in the field of chemical nomenclature led to the prevalent use of line notation. Instead of drawing detailed molecular structures, line notation uses lines, symbols, and characters to convey essential structural information. This is attributed to the computational efficiency with which computers can handle linear data strings. 37 Rapid advancements in computer technology have outpaced the incremental growth of chemical information. The synergy between computer capabilities and chemical knowledge allows for the storage of all existing chemical information on current hardware. The historical challenge of merely storing information in machines has been largely overcome. Current and future efforts are now focused on creating highly efficient systems that deliver chemically relevant information when required. To achieve this, a new chemical language and associated computer programs are being developed, grounded in the innovative information system known as SMILES. 38
@@
SMILES is a line notation used to encode molecular structures, allowing for a linear representation of the molecule so that a computer program can understand the structure. It specifically represents a valence model of a molecule and is widely used in chemistry to represent molecular structures in a compact and human-readable form. The graph of a chemical structure was to be uniquely described, including but not limited to the molecular graph comprising nodes (atoms) and edges (bonds). 39 Since most of the datasets come in SMILES 38 encoded format, they need to be converted into a graph database for use in GNNs. A molecular graph can be represented in the form of $G(n)=(V,E)$ where $n$ is the number of nodes or vertices, i.e the atoms in the molecule and $e_{v,w}$ represents the edges i.e the atomic bonds in the molecule where $e_{v,w}\epsilon E$. A typical example of the molecule in this format has been shown in Figure 1.
Several types of atomic and bond features are present; however, for this study, only the basic features available without data-specific feature engineering are chosen to allow for the model’s generalizability. Based on Figure 1, a sample encoding structure of the features is presented in Tables 2 and 1 along with their encodings. These numerical encodings are then converted to one-hot encoded vectors for use in the geometric dataset. The tables show a small sample of the node and edge feature set.
Table 1: Representation of encoding of edge features
Edge Bond Conjugated Aromatic Is In Ring Connects
Edge 1 Single No No No (1,2)
Edge 2 Double No No No (2,3)
Edge 3 Single No No No (2,4)
@@
Node Atom Type Atomic Number Neighbour Hydrogen Formal Charge
@@
A complete representation of the number of features used for both the node and edge cases has been shown in Tables 3 and 4. They also represent the size of the feature space of the one-hot encoded vectors. Proper dimension size and the type of features with detailed descriptions are presented.
Table 3: Edge feature encoding descriptions
@@
Bond type Single, Double, Triple and Aromatic 4 [0–3]
@@
Table 4: Node feature encoding descriptions
Atom type All possible atoms in the periodic table 117 [0–116]
Bonds Number of bonds the atom is involved in 9 [117–125]
Charge Formal charge of the atom 1 [126]
Hybridization sp, sp2, sp3, sp3d, sp3d2, sp2d, s, other, unspecified 9 [127–135]
Is aromatic Indicates if the atom forms any aromatic system 1 [136]
Is in ring Indicates if the atom forms any ring 1 [137]
Atomic number Atomic number 1 [138]
Hs Number of bonded hydrogen atoms 9 [139–147]